On January 1, 2017, Windsor Company issued 10-year, $2,200,000 face value, 6% bonds, at par. Each $1,000 bond is convertible into 14 shares of Windsor common stock. Windsor’s net income in 2017 was $295,000, and its tax rate was 40%. The company had 97,000 shares of common stock outstanding throughout 2017. None of the bonds were converted in 2017.

(a) Compute diluted earnings per share for 2017. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share

$enter diluted earnings per share rounded to 2 decimal places


(b) Compute diluted earnings per share for 2017, assuming the same facts as above, except that $970,000 of 6% convertible preferred stock was issued instead of the bonds. Each $100 preferred share is convertible into 5 shares of Windsor common stock. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share

$
Answer #1

a) Diluted earning per share = Adjusted net income/Adjusted shares = (295000+79200)/(97000+30800) = 2.93

b) Diluted earning per share = 295000/(97000+48500) = 2.03